# Primordial black holes with anisotropic hair

### v1

@@ IV.2 A toy example
In the early stage of inflation, the energy density of $A$ is negligible, reducing the system to a two-field system. Assuming slow-roll conditions for both $\psi$ and $\phi$, adiabatic and entropy perturbations remain decoupled at early times [49]. Consequently, the curvature perturbation remains conserved on large scales. The ratio $r$ evolves as $r(t)=r_{\text{ini}}[1+2\mathcal{I}(t)^{2}/3]$ due to the anisotropic phase of the late time. We assume that $r_{\text{ini}}\gg 1$. In this regime, $\mathcal{P}_{\mathcal{R}}\simeq\mathcal{P}_{\psi}=(H/\dot{\psi})^{2}|\delta\psi|^{2}=2.1\times 10^{-9}$ dominates the power spectrum on large scales($k\lesssim 1\ \text{Mpc}^{-1}$). At time $N_{f}$ after the time in which the CMB modes exit the horizon, $\mathcal{I}$ becomes significant. Then the power spectrum is dominated by the $\mathcal{P}_{\phi\text{-}A}$ on scale $k_{f}$. One of the examples of this model is shown in FIG. 1. We can obviously observe the anisotropic enhancements and the anisotropic peaks(the black curve).
Figure 1: The power spectrum of $\mathcal{R}$ for our toy example (IV.2). We use the Gaussian profile $\mathcal{I}(N)=\mathcal{I}_{\text{max}}e^{-(N-N_{f})^{2}/(2\Delta N^{2})}$ with $\mathcal{I}_{\text{max}}=9$, $\Delta N=2$, $N_{f}\simeq N_{\text{CMB}}+5$ and $r_{\text{ini}}=10$. The black curve represent peaks $k_{p}(\theta)$ of $\mathcal{P}_{\mathcal{R}}$.
@@ V Conclusion
This mechanism is not restricted to specific models — it can be triggered whenever the choices of $f(\phi)$ and $V(\phi)$ drive inflation into an anisotropic phase. Therefore, once inflation enters an anisotropic phase at a certain time $N_{f}$, it produces a corresponding anisotropic peak in the power spectrum at scale $k_{p}$. However, several aspects require careful consideration. We need to future study their non-Gaussianity(NG) and perturbative control. One may be concerned that the exponential enhancement of perturbations leads to an incredible enhancement on NG. However, it has been shown that the low-energy EFT with imaginary sound speed reduced from strongly non-geodesic attractors of multifield inflation generates only mild NG [32, 31, 33, 34]. Moreover, the perturbative validity is also well controlled in this attractor [31, 32, 35]. We believe that this analysis can be used in our case in this Letter.
The anisotropic features of the peak can be caught by some future observations, such as PBH production. As we know, the PBH production of a statistically anisotropic peak is still absent. It is necessary to figure out the physical details of this case. In particular, anisotropic hair may be imprinted on the initial spin of PBHs, leaving behind intriguing observational signatures [50]. Moreover, enhancements of the curvature power spectrum on small scales are generally accompanied by the production of scalar-induced gravitational waves(SIGWs). This implies that anisotropic features will also manifest as sources in the SIGW spectrum. The frequencies of SIGWs is related to $N_{f}$ as $N_{f}+\log\left(\mathcal{I}\right)\sim\log\left(f_{\text{peak}}/\text{Hz}\right)+37$ [51], which spans a broad spectrum covered by gravitational-wave experiments, such as PTA and LISA. A model-independent discussion of SIGWs sourced from a statistically anisotropic scalar spectrum has been studied in [52, 53]. We provide a specific realization of a statistically anisotropic peak in this Letter. It is interesting to explore signatures of this anisotropic hair in the SIGWs and test it in future observations.

### v2

@@ IV.2 A toy example
In the early stage of inflation, the energy density of $A$ is negligible, reducing the system to a two-field system. Assuming slow-roll conditions for both $\psi$ and $\phi$, adiabatic and entropy perturbations remain decoupled at early times [Gordon:2000hv]. Consequently, the curvature perturbation remains conserved on large scales. The ratio $r$ evolves as $r(t)=r_{\text{ini}}[1+2\mathcal{I}(t)^{2}/3]$ due to the anisotropic phase of the late time. We assume that $r_{\text{ini}}\gg 1$. In this regime, $\mathcal{P}_{\mathcal{R}}\simeq\mathcal{P}_{\psi}=(H/\dot{\psi})^{2}|\delta\psi|^{2}=2.1\times 10^{-9}$ dominates the power spectrum on large scales($k\lesssim 1\ \text{Mpc}^{-1}$). At time $N_{f}$ after the time in which the CMB modes exit the horizon, $\mathcal{I}$ becomes significant. Then the power spectrum is dominated by the $\mathcal{P}_{\phi\text{-}A}$ on scale $k_{f}$. One of the examples of this model is shown in Fig. 1. We can obviously observe the anisotropic enhancements and the anisotropic peaks(the black curve).
Figure 1: The power spectrum of $\mathcal{R}$ for our toy example (IV.2). We use the Gaussian profile $\mathcal{I}(N)=\mathcal{I}_{\text{max}}e^{-(N-N_{f})^{2}/(2\Delta N^{2})}$ with $\mathcal{I}_{\text{max}}=5$, $\Delta N=2$, $N_{f}\simeq N_{\text{CMB}}+5$ and $r_{\text{ini}}=10$. $\mathcal{P}_{\mathcal{A}}=\mathcal{P}_{0}\left|2-4\sin^{2}\theta\right|^{1/2}e^{4\pi\left|\sin{\theta}-\frac{\sqrt{2}}{2}\right|\mathcal{I}}\big|_{N_{v}(k)}$, where $N_{v}(k)$ is given by $p(N_{v})=|m_{\varphi}(N_{v})|$. The black curve represent peaks $k_{p}(\theta)$ of $\mathcal{P}_{\mathcal{R}}$.
@@ V Conclusion
This mechanism is not restricted to specific models—it can be triggered whenever the choices of $f(\phi)$ and $V(\phi)$ drive inflation into an anisotropic phase. Therefore, once inflation enters an anisotropic phase at a certain time $N_{f}$, it produces a corresponding anisotropic peak in the power spectrum at scale $k_{p}$. However, several aspects require careful consideration. We need to future study their non-Gaussianity(NG) and perturbative control. One may be concerned that the exponential enhancement of perturbations leads to an incredible enhancement on NG. However, it has been shown that the low-energy EFT with imaginary sound speed reduced from strongly nongeodesic attractors of multifield inflation generates only mild NG [Bjorkmo:2019qno, Fumagalli:2019noh, Garcia-Saenz:2019njm, Garcia-Saenz:2018vqf]. Moreover, the perturbative validity is also well controlled in this attractor [Fumagalli:2019noh, Bjorkmo:2019qno, Garcia-Saenz:2025jis]. We believe that this analysis can be used in our case in this paper. Moreover, we only consider $\Delta N\gtrsim\log\mathcal{I}$ in this paper. References [Palma:2020ejf, Fumagalli:2020adf] demonstrate that the curvature perturbation power spectrum exhibits oscillatory features when $\Delta N\lesssim\log\mathcal{I}$. It is interesting to explore this case under our anisotropic model. The anisotropic hair may imprint distinctive features in this power spectrum.
The anisotropic features of the peak can be caught by some future observations, such as PBH production. As we know, the PBH production of a statistically anisotropic peak is still absent. It is necessary to figure out the physical details of this case. In particular, anisotropic hair may be imprinted on the initial spin of PBHs, leaving behind intriguing observational signatures [Harada:2024jxl]. Moreover, enhancements of the curvature power spectrum on small scales are generally accompanied by the production of scalar-induced gravitational waves(SIGWs). This implies that anisotropic features will also manifest as sources in the SIGW spectrum. The frequencies of SIGWs is related to $N_{f}$ as $N_{f}+\log\left(\mathcal{I}\right)\sim\log\left(f_{\text{peak}}/\text{Hz}\right)+37$ [Fumagalli:2020nvq], which spans a broad spectrum covered by gravitational-wave experiments, such as PTA and LISA. A model-independent discussion of SIGWs sourced from a statistically anisotropic scalar spectrum has been studied in [Chen:2022qec, Kuang:2023urj]. We provide a specific realization of a statistically anisotropic peak in this paper. Moreover, the scalar modes can be the source of GWs generated during inflation. The GWs are enhanced by the presence of a scalar excited state and overcome the SIGWs [Fumagalli:2021mpc]. It is interesting to explore signatures of this anisotropic hair in the stochastic gravitational wave background(SGWB) and test it in future observations.